# Neutron and $\bm{\gamma}$-ray Discrimination by a Pressurized Helium-4 Based Scintillation Detector

### v1

@@ 1 Introduction
Neutrons as radiation quanta, are found in nature because of their release in various forms of nuclear reactions, most common being the nuclear fission and the $(\alpha,n)$ reactions caused by the natural radioactivity of the remnants of Uranium-Thorium (U-Th) decay chain. Free neutrons are, however, unstable against $\beta$-decay with about 15 minutes of half life. In spite of this fact, free neutrons pass and penetrate through matter since they are charge neutral and deposits partial or full energy through hadronic interactions. These interactions include capture into nuclei and elastic scattering resulting in nuclear recoil of the stopping media. Such interactions take place over a very short time scale compared to the half life of free neutrons, which make it possible to detect them in real life.
Neutrons are produced in large number inside the core of the nuclear reactors and also emerge out of the spent nuclear fuels by spontaneous fission and $(\alpha,n)$ reactions[1]. These neutrons, detected with the help of suitable neutron detectors, are often used to monitor the spent fuel repositories and also at the strategic surveillance stations for monitoring hidden nuclear materials. In that respect, neutron detectors capable of detecting neutrons up to around 10 MeV energy serve the purpose. However, since the neutrons are most often found in a mixed radiation field with dominance of mostly $\gamma$-rays, X-rays and electrons coming out of the same sources, it is important to achieve discrimination between the neutrons and the other radiation quanta before any meaningful result can be extracted.
@@
In recent times, PHe detectors are made commercially available by Arktis Radiation Detectors Ltd, Switzerland[8]. Front-end electronics and data acquisition system is provided as a package for measuring the fast and the thermal neutron fluxes after pulse shape discrimination using Time-over-Threshold (ToT) technique to discriminate between $\gamma$-rays, fast neutrons and the thermal neutrons. Incidentally, the inner wall of the detector is coated with a Lithium-based compound to make the detector sensitive to the thermal neutrons. These neutrons undergo capture by Lithium to produce energetic $\alpha$-particles which result in scintillation inside the detector volume. Corresponding ToT signals are found to be larger than those produced by the fast neutrons. Some technical details about the variant detector model S670, which is not sensitive to the thermal neutrons, can be found elsewhere[9].
@@ 2.2 Measurements with Radioactive sources
The systematic studies were done using different neutron and $\gamma$-ray sources placed at a certain distance from the detector as shown in the photograph (see Fig. 1(a)). Provision for placement of different absorber materials in between for systematic studies was also made. An unmoderated 252Cf spontaneous fission source (half life = 2.645 years, spontaneous fission branching ratio = 3.09%[11]) was used for simultaneous detection of neutrons and $\gamma$-rays. The source emits fast neutrons with average energy of 2.12 MeV. The source was sealed inside a cylindrical PTFE capsule having an opening on one of the flat faces. The aperture diameter was 2 mm, which was sealed with a 100 $\mu$m thick polyethylene terepthalate (PET) window.
@@
An uncalibrated 137Cs monoenergetic (662 keV) $\gamma$-ray source was used for the experiment to identify the $\gamma$-band as distinct from the neutron band. The scatter plot, obtained after exposing the detector with the $\gamma$-ray source, is displayed in the Fig. 3, which shows a single band ($\gamma$-band) as expected.
@@
2.3 Calibration of the detector
Energy calibration of the detector was done using the prompt neutrons from a 252Cf source. It is expected that the energy spectrum of the neutrons originating from the spontaneous fission is Maxwellian. However, after detailed analysis by various groups on the properties of neutron sources, and several meetings under the banner of International Atomic Energy Agency (IAEA), held during 1980 to 1987, the expert group on 252Cf spontaneous fission-based fast neutron sources, had proposed a corrected Maxwellian spectrum, based on experimental data and theoretical estimates[12] over various energy segments from 0.2 MeV to 20 MeV. Accordingly, the corrected Maxwellian spectrum $F(E_{n})$ as function of the kinetic energy ($E_{n}$) of the neutrons for the prompt fission neutrons is given by
A comparison of the experimentally obtained 252Cf prompt neutron spectrum and the one theoretically obtained as above, are shown in the Fig. 5. They are found to be in good agreement over the energy range of  0.7 MeV to  3.2 MeV, which has enabled calibration of the detector. It is important to mention two major points here. a) The upper limit of the available energy range is due to the gain settings of the built-in amplifiers associated with the 4He detector. b) Our comparison, shown in the Fig. 5, manifests that the measured spectrum is constrained by a lower energy cut-off around 0.75 MeV, whereas the estimated spectrum, folded by the efficiency data, has a lower energy cut-off of around 0.35 MeV which arises entirely from the available intrinsic efficiency data in Ref. [14].
There are uncertainties in the calibration which is caused by the sparse data on the intrinsic efficiency. The measured and the simulated efficiency results considerably differ over various energy ranges[14]. Besides, interpolation procedure followed leaves scope for introducing additional uncertainties. Because of these uncertainties, we have adopted a linear calibration without a cut-off. However, a least square fit of the calibration data up to the quadratic term (ie. $E_{n}(x)=a_{0}+a_{1}x+a_{2}x^{2}$, where $x$ is the channel number), results in the following best fit values: $a_{0}=0.00167\pm 0.00212\,{\rm MeV},a_{1}=0.0125\pm 3.93\times 10^{-5}\,{\rm MeV/ch},a_{2}=1.87\times 10^{-8}\pm 1.48\times 10^{-7}\,{\rm MeV/ch^{2}}$. Because of relatively smaller intercept and the quadratic term, our choice of calibration procedure is well-justified. Since identical gain settings were maintained throughout the experiment, we utilized the same energy calibration for the neutrons obtained from different radioactive sources.
It may be noted that the detector is also capable of detecting the $\gamma$-rays due to its electron recoil response, though the intrinsic detection efficiency may be very small compared to that of the neutrons. Though considerable details about the shape of the energy spectral distribution of $\gamma$-rays from spontaneous fission of 252Cf is available[15], similar energy calibration procedure could not be done for the $\gamma$-ray spectra, because of the overlap of the neutron and the $\gamma$-bands (see Fig. 2) at lower energies. However, extending the energy calibration for the neutrons, which essentially relates to the recoil energy of 4He, we refer the common energy scale as equivalent neutron energy $E_{\rm en}$.
@@ 2.4 Measurements with 241Am X-ray source
A 241Am X-ray source, embedded in a glass matrix and sealed inside a cylindrical stainless steel capsule, was also studied with the detector. The source, intended for energy calibration of X-ray detectors, was used in the experiment to determine the response of the detector to low energy photons so that a low energy cut-off for electron recoil could be established. The scatter plot of $P$ vs. $Q_{L}$, obtained after 5 hours of exposure, is shown in the Fig. 6. A clear neutron band, besides the $\gamma$-band, was observed to appear. This was a bit surprising due to the reason that 1) the source was used so far for energy and relative efficiency calibration of the X-ray detectors, and 2) it was never monitored with neutron detectors or dosimeters. Surprisingly, the same exposure time (5 hours) for the 241Am and the 252Cf sources to the PHe detector was used.
@@
In order to confirm that the band is due to the neutrons being detected, attenuation of the band population by high density polyethylene (HDPE) was studied. For this purpose, 4 layers of HDPE, each being 25 mm thick with a total thickness of 100 mm, were placed between the 241Am source and the PHe detector to attenuate the neutrons. The scatter plot ($P$ vs. $Q_{L}$), obtained over the same exposure time, is shown in the Fig. 6. It clearly reveals attenuation of the neutron band. The plots, obtained by gating on the neutron band and projection taken along the $Q_{L}$ axis (see Fig. 7), provide a comparison of the neutron attenuation by the HDPE layers. Low energy neutrons are expected to be attenuated more than those of higher energies, however from the plot, the attenuation factor appears to be fairly independent of energy. This is due to the moderation of the energetic neutrons that enhances counts at low energy and suppresses at higher energy. Similar projection spectrum for the $\gamma$-band is shown in the Fig. 8, where attenuation of $\gamma$-rays is observed. We also have observed build-up of counts in the high energy tail (larger $Q_{L}$). This was investigated in detail through simulation in Sec. 3.4.
@@
Figure 7: (a) Projected spectra on the $Q_{L}$ axis after setting a banana gate on the neutron band with and without the HDPE layers between the source and the detector. Attenuation of neutrons due to HDPE is evident; (b) comparison of the projected neutron spectra recorded after the neutrons travel through different numbers of the HDPE layers.
@@
Figure 8: Same as Fig. 7 except that the banana gate is set on the $\gamma$-band. (a) projected spectra with and without the HDPE layers between the source and the detector. See Sec. 3.4 for the calibration procedure applied in the plot.(b) Comparison of the projected $\gamma$-spectra recorded after the neutrons travel through different numbers of the HDPE layers. The $x$-axis scale on the top displays the equivalent neutron energy $E_{en}$. See Sec. 2.3 for its definition.
@@ 3.2 Neutron Response Systematics by G4
As mentioned in Sec. 1, neutron scatters off a 4He nucleus transferring a part of its kinetic energy ($E_{n}$). The recoiling 4He deposits the recoil energy within the fiducial volume of the detector via ionization, followed by scintillation. The recoiling 4He deposits its entire energy within the volume, which is designated as $E_{dep}$. The $E_{dep}$ spectra for monoenergetic neutrons at different energies ranging from 0.25 to 3 MeV are estimated as shown in the Fig. 10. Since there is considerable spread in the $E_{dep}$ spectra, we have estimated the 90% value of the total area under the spectral profile. Subsequently, we have determined the upper limit of the $E_{dep}$ parameter for the integral under the spectral profile, which matches with the above-mentioned number. We have designated the corresponding upper limit on $E_{dep}$ as the marker ($\Delta E$) of the corresponding recoil signal registered by the detector. A plot of ($\Delta E$) as function of $E_{n}$ is shown in the Fig. 10. The ($\Delta E$) is found to be fairly linear with the incident energy $E_{n}$. Therefore, from the $E_{dep}$ spectrum, an estimate of the incident energy can be done.
@@
Figure 10: (a) Simulated $E_{dep}$ spectrum of the mono-energetic neutrons; (b) $\Delta E$ is plotted against energy of the neutrons.
It may be noted that the recoiling 4He deposits its entire energy within the volume. Also, one neutron is found to cause multiple nuclear recoil during its passage. Therefore, each recoil is also recorded as a distinct event for systematic studies and calculation of S1 signal (to be explained in Sec. 4). The HDPE slabs are cumulatively placed and the simulations are done for 50 million events for each thickness of the absorber. Only the fast neutron events selected by energy cut of $\lesssim 70$ keV, are recorded. Fig. 11 shows the distribution of the $E_{dep}$ corresponding to the fast neutron events, which can be compared with the spectra obtained experimentally (see Fig. 7). Energies of the incident neutrons are obtained from the calibration done with our measurements based on the 252Cf source (see Sec. 2.3). The relative ordering of the plots for different thicknesses of the HDPE absorbers agrees with the experiment. However, because of incomplete information on the scaling between $E_{dep}$ and incident energy of the neutrons, further quantitative comparisons through simulations could not be done.
@@ 3.3 Electron recoil events
$\gamma$-rays falling on the PHe detector cause electrons inside the gas to recoil and thus, result in electron recoil signal. This is demonstrated in our experiment (see Fig. 3) with 137Cs standard $\gamma$-ray source as well. The range of $E_{dep}$ values can be seen to merge with the lower range of $E_{dep}$ values for the neutrons as shown in the Fig. 4, where the spectra for the neutrons and the $\gamma$-rays from 252Cf are overlaid. Applying the energy calibration for the neutrons, we can conclude that the $\gamma$-ray spectrum as seen by the PHe detector extends from $\sim 0.6-1.2$ MeV in neutron equivalent energy $E_{en}$. Reaching to lower energy value in $E_{en}$ is constrained by the threshold of the detector for neutron-$\gamma$ discrimination.
It has already been demonstrated by exposing the detector with 241Am source that the neutrons are attenuated by the HDPE layers (see Sec. 2.4). It is also shown that while the $\gamma$-ray spectrum at the lower channels along the $Q_{L}$ axis get significantly attenuated (see Fig. 8) after passing though the HDPE layers, the higher channels corresponding the larger $E_{dep}$ values get populated as well. For measurements with different number of HDPE layers in between, we observe that after traversing two HDPE layers, the spectrum at the higher channels (i.e, region of interest or RoI $>150$) show a rising trend and a bump around 240 channel number till it reaches saturation at the highest channel. There is not much visible relative change for four HDPE layers. However, the spectral population over the RoI diminish after passing through five HDPE layers. This may be interpreted as due to the production of 2.225 MeV $\gamma$-rays from neutron moderation within the HDPE, followed by absorption in HDPE through the $n+p=d+\gamma$ process. A simulation of the $\gamma$-ray interaction with the detector is done to understand the origin, the nature of the $\gamma$-ray spectra and the underlying systematics of the process.
The $\gamma$ rays falling on the detector will produce electron recoil within the pressurized 4He-volume. The study is more complicated due to the presence of stainless steel housing of the detector which contains high $Z$ elements. Since the $\gamma$-rays are energetic to undergo pair creation, positrons are also produced, which result in the production of 511 keV $\gamma$-rays. Therefore, the electron-recoil events, though largely contain effects due to the electrons, there would be effect due to the positrons as well. Fig. 11 includes the $E_{dep}$ spectra for the electron recoil events, when the neutrons emitted from the 241Am source, are incident on the detector.
@@ 3.4 Gamma-ray Response Systematics by G4
In the simulation procedure, mono-energetic $\gamma$-rays of different energies ($E_{\gamma}$) are allowed to impinge upon the detector. The $E_{dep}$ spectrum of the recoiling electrons is shown in Fig. 12. A peak begins to appear at $E_{dep}\sim 280$ keV as the energy of the incident $\gamma$-ray is increased. The $\gamma$-ray energy threshold for appearance of the peak is found to be around 650 keV. Therefore, the 847 keV $\gamma$-ray, which arises from the ${}^{56}{\rm Fe}(n,n^{\prime}\gamma)$ reaction, would also give rise to the $\sim 280$ keV peak due to the presence of SS housing, when no HDPE slab is placed. A detailed event-by-event tracking reveals that this number corresponds to the average $E_{dep}$ of the recoiling electrons given the geometry (diameter) of the detector. The $\Delta E$ values for the $E_{dep}$ spectra as function of $E_{\gamma}$ is plotted in the Fig. 12, which indicates saturating trend at higher energies. The contour is fitted with a quadratic function, modulated by a smooth step function of the form:
$F(x)=0.5\cdot(1-\tanh((x-p0)/p1))\cdot(p5\cdot x^{2}+p3\cdot x+p2)+0.5\cdot p4\cdot(1+\tanh((x-p0)/p1)),$ (2)
@@
Figure 12: (a) $E_{dep}$ spectrum of mono-energetic $\gamma$-rays, and (b) corresponding 90% contour values are plotted as function of the $\gamma$-ray energy.
The saturating behaviour in the Fig. 12 arises due to the fact that the recoiling electrons do not lose their entire energy within the detector. This can be seen more explicitly in Fig. 13, where the $E_{dep}$ corresponding to the recoiling electrons is plotted against their kinetic energy (KE) for mono-energetic incident $\gamma$-rays of 1 MeV and 2.225 MeV. The kinetic energy of the recoiling electrons increases with $E_{\gamma}$ as expected. The sharp right edge in the plots is the Compton edge. It is evident from Fig. 13 that a significant fraction of the electrons with higher recoil energies lose only $\sim$280 keV of their energy within the detector. Therefore, we can conclude that the peak observed at higher channel numbers in the Fig. 7 corresponds to $\sim 280$ keV of $E_{dep}$.
@@
Figure 13: $E_{dep}$ vs. recoil kinetic energy of ER events from mono-energetic $\gamma$ sources for (a) 1.0 MeV and (b) 2.22 MeV
@@
To further confirm this geometric effect, the simulation was repeated by changing the diameter of the fiducial volume of the detector to 30 mm and 90 mm. Corresponding $E_{dep}$ spectra are plotted in the Fig. 14. A clear shift in the peak can be seen, which is at $\sim$130 keV for 30 mm diameter (solid line) and $\sim$400 keV for 90 mm (dotted line).
@@
Another systematic study was done to understand the effect of the SS cover of the detector. In this case, the SS cover was removed (keeping the original diameter of 60 mm unchanged for the active volume) and the corresponding $E_{dep}$ spectrum was obtained. It was observed that the height of the peak at $\sim$280 keV had reduced in the spectrum indicating that the peak was, indeed, enhanced by the presence of the SS cover. The contribution from pair-production events to the RoI and in the overall spectra was evaluated from the $E_{dep}$ spectrum obtained after vetoing those events where pair-production has taken place. No significant deviation was observed from the spectrum of Fig. 12, indicating that pair-production does not make significant contribution to the spectra. Additionally, it was observed from simulation that in almost all the cases, a single recoil electron is produced from a single $\gamma$-ray. The number of events with two recoil electrons from the incident $\gamma$-ray is almost two orders of magnitude lower than the single recoil events.
@@
Figure 14: $E_{dep}$ spectra with 30 (solid lines), 60 (dot-dashed lines) and 90 mm (dotted lines) diameter of the SS housing.

### v2

@@ 1 Introduction
Neutrons as radiation quanta, are found in nature because of their release in various forms of nuclear reactions, most common being the nuclear fission and the $(\alpha,n)$ reactions caused by the natural radioactivity of the remnants of Uranium-Thorium (U-Th) decay chain. Free neutrons are, however, unstable against $\beta$-decay with about 15 minutes of half life. In spite of this fact, free neutrons pass and penetrate through matter since they are charge neutral and deposits partial or full energy through hadronic interactions. These interactions include capture into nuclei and elastic scattering, resulting in nuclear recoil of the stopping media. Such interactions take place over a very short time scale compared to the half-life of free neutrons, which make it possible to detect them in real life.
Neutrons are produced in large number inside the core of the nuclear reactors and also emerge out of the spent nuclear fuels by spontaneous fission and $(\alpha,n)$ reactions[1]. These neutrons, detected with the help of suitable neutron detectors, are often used to monitor the spent fuel repositories and also at the strategic surveillance stations for monitoring hidden nuclear materials. In that respect, neutron detectors, capable of detecting neutrons up to around 10 MeV energy, serve the purpose. However, since the neutrons are most often found in a mixed radiation field with dominance of mostly $\gamma$-rays, X-rays and electrons coming out of the same sources, it is important to achieve discrimination between the neutrons and the other radiation quanta before any meaningful result can be extracted.
@@
In recent times, PHe detectors are made commercially available by Arktis Radiation Detectors Ltd, Switzerland[8]. Front-end electronics and data acquisition system is provided as a package for measuring the fast and the thermal neutron fluxes after pulse shape discrimination using Time-over-Threshold (ToT) technique to discriminate between $\gamma$-rays, fast neutrons and the thermal neutrons. The inner wall of the detector is coated with a Lithium-based compound to make the detector sensitive to the thermal neutrons. These neutrons undergo capture by Lithium to produce energetic $\alpha$-particles which result in scintillation inside the detector volume. Corresponding ToT signals are found to be larger than those produced by the fast neutrons. Some technical details about the variant detector model S670, which is not sensitive to the thermal neutrons, can be found elsewhere[9].
@@ 2.2 Measurements with Radioactive sources
The systematic studies were done using different neutron and $\gamma$-ray sources placed at a certain distance from the detector as shown in the photograph (see Fig. 1(a)). Provision for placement of different absorber materials in between for systematic studies was also made. An unmoderated 252Cf spontaneous fission source (half-life = 2.645 years, spontaneous fission branching ratio = 3.09%[11]) was used for simultaneous detection of neutrons and $\gamma$-rays. The source emits fast neutrons with average energy of 2.12 MeV. The source, with a strength of $\sim 1-5\,{\rm kBq}$, was sealed inside a cylindrical PTFE capsule having an opening on one of the flat faces. The aperture diameter was 2 mm, which was sealed with a 100 $\mu$m thick polyethylene terepthalate (PET) window.
@@
A 137Cs monoenergetic (662 keV) $\gamma$-ray source, with strength $\sim 30\,{\rm kBq}$, was used for the experiment to identify the $\gamma$-band as distinct from the neutron band. The scatter plot, obtained after exposing the detector with the $\gamma$-ray source, is displayed in the Fig. 3, which shows a single band ($\gamma$-band) as expected.
@@ 2.3 Measurements with 241Am X-ray source
A 241Am X-ray source, embedded in a glass matrix and sealed inside a cylindrical stainless steel capsule, was also studied with the detector. The source, having a strength of $\sim 100\,{\rm MBq}$, is originally intended for energy calibration of X-ray detectors. This source was used in the experiment to determine the response of the detector to low energy photons so that a low energy cut-off for electron recoil could be established. The scatter plot of $P$ vs. $Q_{L}$, obtained after 5 hours of exposure, is shown in the Fig. 5. A clear neutron band, besides the $\gamma$-band, was observed to appear. This was a bit surprising due to the reason that 1) the source was used so far for energy and relative efficiency calibration of the X-ray detectors, and 2) it was never monitored with neutron detectors or dosimeters. Surprisingly, the same exposure time (5 hours) for the 241Am and the 252Cf sources to the PHe detector was used.
In order to confirm that the band is due to the neutrons being detected, attenuation of the band population by high density polyethylene (HDPE) was studied. For this purpose, 4 layers of HDPE, each being 25 mm thick with a total thickness of 100 mm, were placed between the 241Am source and the PHe detector to attenuate the neutrons. The scatter plot ($P$ vs. $Q_{L}$), obtained over the same exposure time, is shown in the Fig. 5.
@@
It clearly reveals attenuation of the neutron band. The plots, obtained by gating on the neutron band and projection taken along the $Q_{L}$ axis (see Fig. 6), provide a comparison of the neutron attenuation by the HDPE layers. Low energy neutrons are expected to be attenuated more than those of higher energies, however from the plot, the attenuation factor appears to be fairly independent of energy. This is due to the moderation of the energetic neutrons that enhances counts at low energy and suppresses at higher energy. Similar projection spectrum for the $\gamma$-band is shown in the Fig. 7, where attenuation of $\gamma$-rays is observed. We also have observed build-up of counts in the high energy tail (larger $Q_{L}$). This was investigated in detail through simulation in Sec. 3.4.
@@
Figure 6: (a) Projected spectra on the $Q_{L}$ axis after setting a banana gate on the neutron band with and without the HDPE layers between the source and the detector. Attenuation of neutrons due to HDPE is evident; (b) comparison of the projected neutron spectra recorded after the neutrons travel through different numbers of the HDPE layers. See Sec. 3.4 for the energy calibration applied here.
@@
Figure 7: Same as Fig. 6 except that the banana gate is set on the $\gamma$-band. (a) projected spectra with and without the HDPE layers between the source and the detector. See Sec. 3.4 for the calibration procedure applied in the plot.(b) Comparison of the projected $\gamma$-spectra recorded after the neutrons travel through different numbers of the HDPE layers. The $x$-axis scale on the bottom displays the electron equivalent energy loss $E_{dep}^{ee}$ of neutron. See Sec. 3.3 for its definition.
@@ 3.2 Neutron Response Systematics by G4
As mentioned in Sec. 1, neutron scatters off a 4He nucleus transferring a part of its kinetic energy ($E_{n}$). The recoiling 4He deposits the recoil energy within the fiducial volume of the detector via ionization, followed by scintillation. The recoiling 4He deposits its entire energy within the volume, which is designated as $E_{dep}$. The $E_{dep}$ spectra for monoenergetic neutrons at different energies ranging from 0.25 to 3 MeV are estimated as shown in the Fig. 9. Since there is spread in the $E_{dep}$ spectra, we have estimated the 90% value of the total area under the spectral profile. Subsequently, we have determined the upper limit of the $E_{dep}$ parameter for the integral under the spectral profile, which matches with the above-mentioned number. We have designated the corresponding upper limit on $E_{dep}$ as the marker ($\Delta E$) of the corresponding recoil signal registered by the detector. A plot of ($\Delta E$) as function of $E_{n}$ is shown in the Fig. 9. The ($\Delta E$) is found to be fairly linear with the incident energy $E_{n}$. Therefore, from the $E_{dep}$ spectrum, an estimate of the incident energy may be done.
@@
Figure 9: (a) Simulated $E_{dep}$ spectrum of the mono-energetic neutrons; (b) $\Delta E$ is plotted against energy of the neutrons. See Sec.3.2 for the definition of the $\Delta E$ parameter.
It may be noted that the recoiling 4He deposits its entire energy within the volume. Also, one neutron is found to cause multiple nuclear recoil during its passage. Therefore, each recoil is also recorded as a distinct event for systematic studies and calculation of S1 signal (to be explained in Sec. 4). The HDPE slabs are cumulatively placed and the simulations are done for 50 million events for each thickness of the absorber. Only the fast neutron events selected by energy cut of $\lesssim 70$ keV, are recorded. Fig. 10 shows the distribution of the $E_{dep}$ corresponding to the fast neutron events, which can be compared with the spectra obtained experimentally (see Fig. 6). The relative ordering of the plots for different thicknesses of the HDPE absorbers agrees with the experiment.
@@ 3.3 Electron recoil events
$\gamma$-rays falling on the PHe detector cause electrons inside the gas to recoil and thus, result in electron recoil signal. This is demonstrated in our experiment (see Fig. 3) with 137Cs standard $\gamma$-ray source as well. The range of $E_{dep}$ values can be seen to merge with the lower range of $E_{dep}$ values for the neutrons as shown in the Fig. 4, where the spectra for the neutrons and the $\gamma$-rays from 252Cf are overlaid.
It has already been demonstrated by exposing the detector with 241Am source that the neutrons are attenuated by the HDPE layers (see Sec. 2.3). It is also shown that while the $\gamma$-ray spectrum at the lower channels along the $x$-axis get significantly attenuated (see Fig. 7) after passing though the HDPE layers, the higher channels corresponding the larger $E_{dep}$ values get populated as well. For measurements with different number of HDPE layers in between, we observe that after traversing two HDPE layers, the spectrum at the higher channels (i.e, region of interest or RoI $>150$) show a rising trend and a bump around 240 channel number till it reaches saturation at the highest channel. There is not much visible relative change for four HDPE layers. However, the spectral population over the RoI diminish after passing through five HDPE layers. This may be interpreted as due to the production of 2.225 MeV $\gamma$-rays from neutron moderation within the HDPE, followed by absorption in HDPE through the $n+p=d+\gamma$ process. A simulation of the $\gamma$-ray interaction with the detector is done to understand the origin, the nature of the $\gamma$-ray spectra and the underlying systematics of the process.
The $\gamma$ rays falling on the detector will produce electron recoil within the pressurized 4He-volume. The study is more complicated due to the presence of stainless steel housing of the detector which contains high $Z$ elements. Since the $\gamma$-rays are energetic to undergo pair creation, positrons are also produced, which result in the production of 511 keV $\gamma$-rays. Therefore, the electron-recoil events, though largely contain effects due to the electrons, there would be effect due to the positrons as well. Fig. 10 includes the energy loss spectra for the electron recoil events, when the neutrons emitted from the 241Am source, are incident on the detector. Henceforth, we will designate the energy loss by the electrons as $E_{dep}^{ee}$ to distinguish it from $E_{dep}$ due to the neutrons.
@@ 3.4 Gamma-ray Response Systematics by G4
In the simulation procedure, mono-energetic $\gamma$-rays of different energies ($E_{\gamma}$) are allowed to impinge upon the detector. The $E_{dep}^{ee}$ spectrum of the recoiling electrons is shown in Fig. 11. A peak begins to appear at $E_{dep}^{ee}\sim 280$ keV as the energy of the incident $\gamma$-ray is increased. The $\gamma$-ray energy threshold for appearance of the peak is found to be around 650 keV. Therefore, the 847 keV $\gamma$-ray, which arises from the ${}^{56}{\rm Fe}(n,n^{\prime}\gamma)$ reaction, would also give rise to the $\sim 280$ keV peak due to the presence of SS housing, when no HDPE slab is placed. A detailed event-by-event tracking reveals that this number corresponds to the most probable $E_{dep}^{ee}$ of the recoiling electrons given the geometry (diameter) of the detector. The $\Delta E$ values for the $E_{dep}^{ee}$ spectra as function of $E_{\gamma}$ is plotted in the Fig. 11, which indicates saturating trend at higher energies. The contour is fitted with a quadratic function, modulated by a smooth step function of the form:
$F(x)=0.5\cdot(1-\tanh((x-p0)/p1))\cdot(p5\cdot x^{2}+p3\cdot x+p2)+0.5\cdot p4\cdot(1+\tanh((x-p0)/p1))$
@@
Figure 11: (a) $E_{dep}^{ee}$ spectrum of mono-energetic $\gamma$-rays, and (b) corresponding 90% contour values are plotted as function of the $\gamma$-ray energy.
The saturating behaviour in the Fig. 11 arises due to the fact that the recoiling electrons do not lose their entire energy within the detector. This can be seen more explicitly in Fig. 12, where the $E_{dep}^{ee}$ is plotted against their kinetic energy (KE) for mono-energetic incident $\gamma$-rays of 1 MeV and 2.225 MeV. The kinetic energy of the recoiling electrons increases with $E_{\gamma}$ as expected. The sharp right edge in the plots is the Compton edge. It is evident from Fig. 12 that a significant fraction of the electrons with higher recoil energies lose only $\sim$280 keV of their energy within the detector. Therefore, we can conclude that the peak observed at higher channel numbers in the Fig. 6 corresponds to $\sim 280$ keV of $E_{dep}^{ee}$.
@@
Figure 12: $E_{dep}^{ee}$ vs. recoil kinetic energy of ER events from mono-energetic $\gamma$ sources for (a) 1.0 MeV and (b) 2.22 MeV
@@
To further confirm this geometric effect, the simulation was repeated by changing the diameter of the fiducial volume of the detector to 30 mm and 90 mm. Corresponding $E_{dep}^{ee}$ spectra are plotted in the Fig. 13. A clear shift in the peak can be seen, which is at $\sim$130 keV for 30 mm diameter (solid line) and $\sim$450 keV for 90 mm (dotted line).
@@
Another systematic study was done to understand the effect of the SS cover of the detector. In this case, the SS cover was removed (keeping the original diameter of 60 mm unchanged for the active volume) and the corresponding $E_{dep}^{ee}$ spectrum was obtained. It was observed that the height of the peak at $\sim$280 keV had reduced in the spectrum indicating that the peak was, indeed, enhanced by the presence of the SS cover. The contribution from pair-production events to the RoI and in the overall spectra was evaluated from the $E_{dep}^{ee}$ spectrum obtained after vetoing those events where pair-production has taken place. No significant deviation was observed from the spectrum of Fig. 11, indicating that pair-production does not make significant contribution to the spectra. Additionally, it was observed from simulation that in almost all the cases, a single recoil electron is produced from a single $\gamma$-ray. The number of events with two recoil electrons from the incident $\gamma$-ray is almost two orders of magnitude lower than the single recoil events.
@@
Figure 13: $E_{dep}^{ee}$ spectra with 30 (solid lines), 60 (dot-dashed lines) and 90 mm (dotted lines) diameter of the SS housing.